3x -y =-7
x + y = -9
ok I think I need to rewrite the equation with y= 3(x-x)-7

Answer by rapaljer(4551)

Just add the two equations together:

3x -y =-7
x + y = -9
-----------
4x = -16
x=-4

Substitute the x=-4 back into either equation. The second equation will be easier:

x+y=-9
-4 +y = -9
Add +4 to each side:
-4+y+4=-9+4
y=-5

The solution is the point x=-4, y=-5 or (-4,-5).

R^2 Retired from SCC